# Relaxation of first-class constraints and the quantization of gauge theories: from “matter without matter” to the reappearance of time in quantum gravity

### v1

@@ III.7.1 Canonical general relativity
The first-class algebra of general relativity is given by the Poisson brackets (suppressing the temporal dependence for brevity)
@@ III.7.2 Fock–Stueckelberg approach to general relativity
From the variation of the Fock–Stueckelberg action (26), we obtain the Hilbert energy-momentum tensor of the effective matter, [^47] In the functional variation $\delta\mathcal{F}=\varepsilon\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\mathcal{F}(t,\mathbf{x};g_{\mu\nu}+\varepsilon\eta_{\mu\nu}]=:\varepsilon\int\mathrm{d}^{3}x\ \frac{\delta\mathcal{F}}{\delta g_{\mu\nu}}\eta_{\mu\nu}(t,\mathbf{x})$ (cf. Footnote 3), as the metric and its (arbitrary) variation $\eta_{\mu\nu}$ are symmetric tensors, we must symmetrize the functional derivative $\delta\mathcal{F}/\delta g_{\mu\nu}$ with respect to the $\mu,\nu$ indices. For example, from Eq. (93), we find $\delta g^{\mu\nu}(\tau,\mathbf{x})/\delta g_{\rho\sigma}(\tau,\mathbf{y})=-\delta(\mathbf{x},\mathbf{y})(g^{\mu\rho}g^{\sigma\nu}+g^{\mu\sigma}g^{\rho\nu})_{(\tau,\mathbf{x})}/2$.
@@
What can we conclude about the function $\Lambda(h)$ that defines the gauge fixing? And what about the new multiplier field $w$ that characterizes the effective fluid? First, notice that the constraints $\chi\approx 0,\phi_{\perp}+w\approx 0,\phi_{a}\approx 0$ are the ‘primary’ constraints in the Fock–Stueckelberg theory, as they follow from the Fock–Stueckelberg action itself. The constraints $\chi\approx 0$ and $w\approx-\phi_{\perp}$ fix the multipliers $N$ and $w$, respectively, at all times, and their preservation in time fixes the corresponding derivatives of the multipliers. On the other hand, we must proceed with the analysis of the constraint conservation (à la Rosenfeld–Bergmann–Dirac Rosenfeld ; SalisburyRosenfeld ; Dirac3 ; RBD ; HT:book ; Bergmann ), especifically with respect to the (“unaltered”) spatial diffeomorphism constraints $\phi_{a}$ to determine whether new (‘secondary’) constraints appear in the theory and whether the Fock–Stueckelberg theory defined from the parent-theory gauge choice given in Eq. (91) is consistent. We can use the field equations (30) together with the first-class algebra given in Eqs. (90) to find
@@
In this way, the application of the Fock–Stückelberg approach to general relativity naturally leads to new degrees of freedom [coming from the $w$ multiplier(s)] that can be effectively identified with different kinds of matter, whose form depends on the chosen gauge condition(s). This effective matter appears in addition to the matter content that was already present in the parent theory. If the parent theory is purely gravitational, the corresponding Fock–Stueckelberg theory will feature a gravitational field and an effective fluid. Similarly to Dirac’s electrodynamics, we thus have “matter without matter” in the Fock–Stueckelberg theory.
In Kaplan1 , this emergence of effective matter from a relaxation of the first-class constraints in general relativity was rediscovered. It is also important to mention different approaches across the literature, which yield similar results or employ a similar reasoning to the Fock–Stueckelberg approach. The treatment of constrained Hamiltonian systems via the “unfree gauge symmetry” formalism was discussed in the series of papers unfree ; unfree1 ; unfree2 ; unfree3 . The essential idea is to consider gauge transformations where $\varepsilon$ (cf. Sec. II.1) are restricted by additional conditions, namely, they must satisfy a set of partial differential equations. This restriction of the gauge transformations leads to a new theory, just as in the case of the Fock–Stueckelberg method. Indeed, the simplest example is again unimodular gravity, where the corresponding equations are simply the conditions of the transverseness, $\nabla_{\mu}\varepsilon^{\mu}=0$ (with $\mu$ a spacetime index). Moreover, unimodular gravity was analyzed from the point of view of restricted gauge theories in Barv-Nest . Finally, in the so-called TDiff theories (see, e.g., Tdiff ), the matter part of the action is invariant only with respect to the transverse diffeomorphisms, which leads to a weaker version of unimodular gravity. Some models of similar theories were considered in Tdiff1 , where the details of the breaking of diffeomorphism invariance were discussed. The reader is also referred to Percacci for a rather broad analysis of gravity theories with different sets of constraints or different structures of gauge algebras. All in all, the Fock–Stueckelberg approach can be seen as one of the oldest and simplest of these approaches where the gauge symmetry is modified to yield a new theory from a parent.
@@ III.7.3 Homogeneous cosmology
where $N\approx\Lambda$. Notice that the terms proportional to $w_{0}$ constitute a modification of the standard Friedmann equations, which can indeed be interpreted as a constant equation-of-state parameter $\mathscr{W}$ of an effective fluid. Thus, in particular, by assuming $w_{0}/\Lambda_{0}>0$, we have: effective dust (dark matter) for $\mathscr{W}=0$, for which we recover the gauge $N\approx\Lambda_{0}$; [^51] A generalized unimodular gravity with effective dust matter and $N=1$ gauge was also considered in the book Burlan and the series of articles cited therein. an effecitve cosmological constant (dark energy) for $\mathscr{W}=-1$ (which corresponds to the unimodular gauge condition in the parent theory); an effective curvature contribution for $\mathscr{W}=-1/3$; and effective radiation contribution for $\mathscr{W}=1/3$.
@@ III.7.4 Time from constants of Nature
for any phase-space function $f\equiv f(a,\varphi)$, and the Fock–Stueckelberg Hamiltonian is conserved but not necessarily zero (the constraint is relaxed),
@@
with $f\equiv f(a,\varphi)$ being a phase-space function. As $\mathscr{E}_{0}$ is a conserved quantity (constant of Nature) in this toy model, we are thus in a similar situation to the relativistic particle case discussed in Sec. III.5.2, with $\mathscr{E}_{0}$ being analogous to the particle’s rest energy. In this way, we can promote $p_{\eta}\coloneqq\Lambda_{0}\mathscr{E}_{0}$ to a new, conserved degree of freedom, and we enlarge the set of variables from $(a,p_{a},\varphi,p_{\varphi})$ to include the new canonical pair $(\eta,p_{\eta})$ by adding to the action the new term $p_{\eta}\dot{\eta}$. As in the particle case, $\eta$ is a cyclic variable that serves a physical clock that defines time as a canonical conjugate to a constant of Nature. By imposing the time parametrization $\eta(\tau)=\tau$, we fix the lapse function to be $N=\Lambda_{0}a^{3\mathscr{W}}$, and we can solve the constraint $a^{3\mathscr{W}}\phi_{\perp}+\mathscr{E}_{0}\approx 0$ for $\mathscr{E}_{0}$ to obtain the on-shell gauge-fixed equations of motion
@@ IV The reappearance of time in quantum gravity
which imply a timeless constraint $\hat{H}\ket{\Psi}=\lambda^{a}(\tau)\hat{\phi}_{a}\ket{\Psi}=0$ (spatial integration and summation over discrete indices implied). Due to the structure of ϕsubscriptitalic-$\phi_{\perp}$ (cf. Sec. III.7.1), the timeless quantum constraints lead to a generalization of the Klein–Gordon equation for the relativistic particle, and thus the problem of time in quantum gravity is inevitably linked to difficulties with the definition of the Hilbert space. [^54] See, however, Marolf ; Giulini for way a promising way to define a positive-definite inner product at least for symmetry-reduced toy models. In fact, the Hamiltonian constraint of canonical quantum cosmology of homogeneous models coincides with that obtained from mechanical models discussed in Sec. III.5, which is nothing but a Klein–Gordon equation for a relativistic particle with a general potential and a general background (configuration-space) metric.
@@
due to the fact that the time translations are no longer gauged. One can thus speak of the “reappearance of time,” which is nothing but the breaking of spacetime diffeomorphism invariance engendered by the Fock–Stueckelberg construction. It is important to emphasize again that this construction is not related to the consistency of the quantization procedure itself, but rather to the definition of a different classical theory to be subsequently quantized. It makes the quantization more conceptually and technically straightforward at the cost of the theory’s local symmetry, and it was one of the motivations for Stueckelberg to develop his approach for the relativistic particle (by breaking the worldline diffeomorphism invariance), as well as for modern researchers. Indeed, from the “original” unimodular gravity to its generalizations HT-CC-1 ; UnruhUni ; UnruhWald ; HT-CC-2 ; Kaplan1 ; we ; Barv ; Barv1 ; Kam-Tron-Ven , and with several proposals to define time as the canonical conjugate to different constants of Nature CG1 ; CG2 ; CG3 ; GT1 ; GT2 ; GT3 ; GT4 ; Magueijo1 ; Magueijo2 ; Magueijo3 ; Magueijo4 ; Magueijo5 ; Magueijo6 ; Magueijo7 ; Kaplan , this strategy has led to a series of articles claiming to solve the problem of time in quantum gravity.
@@ V Conclusions
Generally, one can say that in a theory with gauge symmetries or in a theory with first-class constraints, the relaxation of some of these constraints implies the appearance of new physical degrees of freedom. This is the essence of the Fock–Stueckelberg approach that we have discussed in this article. The idea is rather simple, and our goal was to present a comprehensive overview of this method, which is often independently rediscovered in the literature, and many sources do not include a discussion of the origins and physical meaning of the formalism.
@@
It is clear that the Fock–Stueckelberg approach generates a classical theory, possibly different from its parent theory. The question of quantization is secondary (the Fock–Stueckelberg theory can be subsequently quantized with the usual methods, from Dirac to BRST quantization Hanson ; HT:book ). This goes against some recent claims made in Kaplan ; Kaplan1 ; Kaplan2 that fixing the gauge à la Fock and Stueckelberg (i.e., fixing the gauge at the level of the action before variation) is necessary in order to make the quantum theory or, more precisely, the quantum Hamiltonian and the path integral of gauge theories well defined. Those can be well defined, presumably, via the usual methods, provided field-theoretic divergences (which come from functional derivatives, for example) can be regularized, an issue about which the Fock–Stueckelberg approach remains silent. Thus, the physical meaning of this formalism need not be tied to the consistency of the quantization of gauge theories, and the fundamental quantum theory need not break gauge invariance. However, it is noteworthy that, since the Fock–Stueckelberg mechanism often leads to a more straightforward quantization, it is often motivated with that end in mind since the original works of Fock and Stueckelberg Fock ; Stueck ; Stueck1 ; Stueck2 .
@@ Acknowledgements.
The work of R.C., A.K. and A.T. was partially supported by the INFN Grant FLAG. The work by F.G.P. was partially supported by the INFN Grant STEFI. The work of L.C. is supported by the Basque Government Grant IT1628-22, and by the Grant PID2021-123226NB-I00 (funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”). It is also partly funded by the IKUR 2030 Strategy of the Basque Government.

### v2

@@ III.7.1 Canonical general relativity
The first-class algebra of general relativity is given by the Poisson brackets (suppressing the temporal and field dependence for brevity)
@@ III.7.2 Fock–Stueckelberg approach to general relativity
From the variation of the Fock–Stueckelberg action (26), we obtain the Hilbert energy-momentum tensor of the effective matter, [^47] In the functional variation $\delta\mathcal{F}=\varepsilon\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\mathcal{F}(\tau,\mathbf{x};g_{\mu\nu}+\varepsilon\eta_{\mu\nu}]=:\varepsilon\int\mathrm{d}^{3}x\ \frac{\delta\mathcal{F}}{\delta g_{\mu\nu}}\eta_{\mu\nu}(\tau,\mathbf{x})$ (cf. Footnote 3), as the metric and its (arbitrary) variation $\eta_{\mu\nu}$ are symmetric tensors, we must symmetrize the functional derivative $\delta\mathcal{F}/\delta g_{\mu\nu}$ with respect to the $\mu,\nu$ indices. For example, from Eq. (93), we find $\delta g^{\mu\nu}(\tau,\mathbf{x})/\delta g_{\rho\sigma}(\tau,\mathbf{y})=-\delta(\mathbf{x},\mathbf{y})(g^{\mu\rho}g^{\sigma\nu}+g^{\mu\sigma}g^{\rho\nu})_{(\tau,\mathbf{x})}/2$.
@@
What can we conclude about the function $\Lambda(h)$ that defines the gauge fixing? And what about the new multiplier field $w$ that characterizes the effective fluid? First, notice that the constraints $\chi\approx 0,\phi_{\perp}+w\approx 0,\phi_{a}\approx 0$ are the ‘primary’ constraints in the Fock–Stueckelberg theory, as they follow from the Fock–Stueckelberg action itself. The constraints $\chi\approx 0$ and $w\approx-\phi_{\perp}$ fix the multipliers $N$ and $w$, respectively, at all times, and their preservation in time fixes the corresponding derivatives of the multipliers. On the other hand, we must proceed with the analysis of the constraint conservation (à la Rosenfeld–Bergmann–Dirac Rosenfeld ; SalisburyRosenfeld ; Dirac3 ; RBD ; HT:book ; Bergmann ), specifically with respect to the (“unaltered”) spatial diffeomorphism constraints $\phi_{a}$ to determine whether new (‘secondary’) constraints appear in the theory and whether the Fock–Stueckelberg theory defined from the parent-theory gauge choice given in Eq. (91) is consistent. We can use the field equations (30) together with the first-class algebra given in Eqs. (90) to find (suppressing the dependence of the constraints on the fields)
@@
In this way, the application of the Fock–Stueckelberg approach to general relativity naturally leads to new degrees of freedom [coming from the $w$ multiplier(s)] that can be effectively identified with different kinds of matter, whose form depends on the chosen gauge condition(s). This effective matter appears in addition to the matter content that was already present in the parent theory. If the parent theory is purely gravitational, the corresponding Fock–Stueckelberg theory will feature a gravitational field and an effective fluid. Similarly to Dirac’s electrodynamics, we thus have “matter without matter” in the Fock–Stueckelberg theory.
In Kaplan1 , this emergence of effective matter from a relaxation of the first-class constraints in general relativity was rediscovered. It is also important to mention different approaches across the literature, which yield similar results or employ a similar reasoning to the Fock–Stueckelberg approach. The treatment of constrained Hamiltonian systems via the “unfree gauge symmetry” formalism was discussed in the series of papers unfree ; unfree1 ; unfree2 ; unfree3 . The essential idea is to consider gauge transformations where $\varepsilon$ (cf. Sec. II.1) are restricted by additional conditions, namely, they must satisfy a set of partial differential equations. This restriction of the gauge transformations leads to a new theory, just as in the case of the Fock–Stueckelberg method. Indeed, the simplest example is again unimodular gravity, where the corresponding equations are simply the conditions for transverseness, $\nabla_{\mu}\varepsilon^{\mu}=0$ (with $\mu$ a spacetime index). Moreover, unimodular gravity was analyzed from the point of view of restricted gauge theories in Barv-Nest . Finally, in the so-called TDiff theories (see, e.g., Tdiff ), the matter part of the action is invariant only with respect to the transverse diffeomorphisms, which leads to a weaker version of unimodular gravity. Some models of similar theories were considered in Tdiff1 , where the details of the breaking of diffeomorphism invariance were discussed. The reader is also referred to Percacci for a rather broad analysis of gravity theories with different sets of constraints or different structures of gauge algebras. All in all, the Fock–Stueckelberg approach can be seen as one of the oldest and simplest of these approaches where the gauge symmetry is modified to yield a new theory from a parent.
@@ III.7.3 Homogeneous cosmology
where $N\approx\Lambda$. Notice that the terms proportional to $w_{0}$ constitute a modification of the standard Friedmann equations, which can indeed be interpreted as contributions of an effective fluid with a constant equation-of-state parameter $\mathscr{W}$. Thus, in particular, by assuming $w_{0}/\Lambda_{0}>0$, we have: effective dust (dark matter) for $\mathscr{W}=0$, for which we recover the gauge $N\approx\Lambda_{0}$; [^51] A generalized unimodular gravity with effective dust matter and $N=1$ gauge was also considered in the book Burlan and the series of articles cited therein. an effective cosmological constant (dark energy) for $\mathscr{W}=-1$ (which corresponds to the unimodular gauge condition in the parent theory); an effective curvature contribution for $\mathscr{W}=-1/3$; and effective radiation contribution for $\mathscr{W}=1/3$.
@@ III.7.4 Time from constants of Nature
for any phase-space function $f\equiv f(a,p_{a};\varphi,p_{\varphi})$, and the Fock–Stueckelberg Hamiltonian is conserved but not necessarily zero (the constraint is relaxed),
@@
with $f\equiv f(a,p_{a};\varphi,p_{\varphi})$ being a phase-space function. As $\mathscr{E}_{0}$ is a conserved quantity (constant of Nature) in this toy model, we are thus in a similar situation to the relativistic particle case discussed in Sec. III.5.2, with $\mathscr{E}_{0}$ being analogous to the particle’s rest energy. In this way, we can promote $p_{\eta}\coloneqq\Lambda_{0}\mathscr{E}_{0}$ to a new, conserved degree of freedom, and we enlarge the set of variables from $(a,p_{a},\varphi,p_{\varphi})$ to include the new canonical pair $(\eta,p_{\eta})$ by adding to the action the new term $p_{\eta}\dot{\eta}$. As in the particle case, $\eta$ is a cyclic variable that serves a physical clock that defines time as a canonical conjugate to a constant of Nature. By imposing the time parametrization $\eta(\tau)=\tau$, we fix the lapse function to be $N=\Lambda_{0}a^{3\mathscr{W}}$, and we can solve the constraint $a^{3\mathscr{W}}\phi_{\perp}+\mathscr{E}_{0}\approx 0$ for $\mathscr{E}_{0}$ to obtain the on-shell gauge-fixed equations of motion
@@ IV The reappearance of time in quantum gravity
which imply a timeless constraint $\hat{H}\ket{\Psi}=\lambda^{a}(\tau)\hat{\phi}_{a}\ket{\Psi}=0$ (spatial integration and summation over discrete indices implied). Due to the structure of ϕsubscriptitalic-$\phi_{\perp}$ (cf. Sec. III.7.1), the timeless quantum constraints lead to a generalization of the Klein–Gordon equation for the relativistic particle, and thus the problem of time in quantum gravity is inevitably linked to difficulties with the definition of the Hilbert space. [^54] See, however, Marolf ; Giulini for a promising way to define a positive-definite inner product at least for symmetry-reduced toy models. In fact, the Hamiltonian constraint of canonical quantum cosmology of homogeneous models coincides with that obtained from mechanical models discussed in Sec. III.5, which is nothing but a Klein–Gordon equation for a relativistic particle with a general potential and a general background (configuration-space) metric.
@@
due to the fact that the time translations are no longer gauged. One can thus speak of the “reappearance of time,” which is nothing but the breaking of spacetime diffeomorphism invariance engendered by the Fock–Stueckelberg construction. It is important to emphasize again that this construction is not related to the consistency of the quantization procedure itself, but rather to the definition of a different classical theory to be subsequently quantized. It makes the quantization more conceptually and technically straightforward at the cost of the theory’s local symmetry, and it was one of the motivations for Stueckelberg to develop his approach for the relativistic particle (by breaking the worldline diffeomorphism invariance), as well as for modern researchers. Indeed, from the “original” unimodular gravity to its generalizations HT-CC-1 ; UnruhUni ; UnruhWald ; HT-CC-2 ; Kaplan1 ; we ; Barv ; Barv1 ; Kam-Tron-Ven ; Kara , and with several proposals to define time as the canonical conjugate to different constants of Nature CG1 ; CG2 ; CG3 ; GT1 ; GT2 ; GT3 ; GT4 ; Magueijo1 ; Magueijo2 ; Magueijo3 ; Magueijo4 ; Magueijo5 ; Magueijo6 ; Magueijo7 ; Kaplan , this strategy has led to a series of articles claiming to solve the problem of time in quantum gravity.
@@ V Conclusions
Generally, one can say that in a theory with gauge symmetries or in a theory with first-class constraints, the relaxation of some of these constraints implies the appearance of new physical degrees of freedom. This is the essence of the Fock–Stueckelberg approach that we have discussed in this article. The idea is rather simple, and it is often independently rediscovered in the literature, whereas its origins and physical meaning frequently remain unclear. Our goal was to present a comprehensive overview of a method that generalizes the original works of Fock and Stueckelberg, which we refer to as the Fock–Stueckelberg theory.
@@
In the literature, the many methods that are related to the Fock–Stueckelberg approach may thus lead to useful models of the sources of the relaxed gauge theories. In this regard, it is worthwhile to mention some recent works that have appeared after the submission of the present article. In DelGrosso , the simultaneous relaxation of the constraints of general relativity and of electromagnetism (a Dirac-type electrodynamics) produces a very particular kind of electric charges. Moreover, it also possible that new developments may lead to extensions of the Fock–Stueckelberg mechanism. In MagueijoNovo ; MagueijoNovo2 , preferred types of foliations (in particular, geodesic foliations) are considered in a type of relaxation of general relativity. There, it is shown that a kind of “dark dust” and more general types of matter may emerge. Interestingly, one seems to obtain results that are different from the applications of generalized unimodular gravity for effective matter fields other than “dark dust.” This is because the effective fluid has anisotropic stress under certain conditions. [^56] We thank João Magueijo for bringing these results to our attention. The particular case of “dark dust” had also been considered in the older reference Mukohyama (see also Kobakhidze ), without, however, a discussion of the connection with (generalized) unimodular gravity or the Fock–Stueckelberg gauge fixation, but including the discussion of another formalism for relaxation of the Hamiltonian constraint: Hořava–Lifschitz gravity. It would be interesting to ascertain whether the recent results of MagueijoNovo ; MagueijoNovo2 are equivalent to a Fock–Stueckelberg gauge fixation or to an extension of it.
@@
It is clear that the Fock–Stueckelberg approach generates a classical theory, possibly different from its parent theory. The Fock–Stueckelberg theory can be subsequently quantized with the usual methods, from Dirac to BRST quantization Hanson ; HT:book . This goes against some recent claims made in Kaplan ; Kaplan1 ; Kaplan2 that fixing the gauge à la Fock and Stueckelberg (i.e., fixing the gauge at the level of the action before variation) is necessary in order to make the quantum theory or, more precisely, the quantum Hamiltonian and the path integral of gauge theories well defined. Those can be well defined, presumably, via the usual methods, provided field-theoretic divergences (which come from functional derivatives, for example) can be regularized, an issue about which the Fock–Stueckelberg approach remains silent. Thus, the physical meaning of this formalism need not be tied to the consistency of the quantization of gauge theories, and the fundamental quantum theory need not break gauge invariance. However, it is noteworthy that, since the Fock–Stueckelberg mechanism often leads to a more straightforward quantization, it is often motivated with that end in mind since the original works of Fock and Stueckelberg Fock ; Stueck ; Stueck1 ; Stueck2 .
@@ Acknowledgements.
The authors thank the anonymous referees for their constructive feedback. The work of R.C., A.K. and A.T. was partially supported by the INFN Grant FLAG. The work by F.G.P. was partially supported by the INFN Grant STEFI. The work of L.C. is supported by the Basque Government Grant IT1628-22, and by the Grant PID2021-123226NB-I00 (funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”). It is also partly funded by the IKUR 2030 Strategy of the Basque Government.
Appendix A Noether charge
In this appendix, we recall the derivation of the equation for the Noether charge. The reader is referred to HT:book ; Pons ; Sundermeyer ; Chataig:Thesis ; Hanson ; SalisburyRosenfeld for further details.
In $d+1$ spacetime dimensions, let us consider the Lagrangian
$L=\int\mathrm{d}^{d}x\,\mathcal{L}(\xi(\tau,\mathbf{x}),\partial_{\mu}{\xi}(\tau,\mathbf{x}))\ ,$ (118)
for the fields $\xi^{i}(\tau,\mathbf{x})$ ($i=1,\ldots,X$), with $(\tau,\mathbf{x})$ comprising the spacetime point $x^{\mu}$ and $\partial_{\mu}\xi\equiv\partial\xi/\partial x^{\mu}$. We consider field derivatives only up to first order for simplicity. In the mechanical case, $d=0$ and there is no space nor spatial derivatives nor spatial integration. Now consider the one-parameter family of transformations
$x^{\mu}$ $\mapsto x^{\mu}(s)\ ,$ (119)
$\xi(\tau,\mathbf{x})$ $\mapsto\xi(\tau(s),\mathbf{x}(s);s)\ ,$
which are assumed to be continuously connected to the identity, and they may include not only changes in the $\xi$ fields themselves but also in the spacetime point (with $x^{\mu}(0)=x^{\mu}$). We have
$\frac{\mathrm{d}}{\mathrm{d}s}\xi(\tau(s),\mathbf{x}(s);s)=\frac{\partial\xi}{\partial s}+\frac{\partial x^{\mu}}{\partial s}\frac{\partial\xi}{\partial x^{\mu}}\ ,$ (120)
where we include a tacit summation over repeated indices, as usual, and we note that $[\partial_{s},\partial_{\mu}]=0$. The partial derivative $\partial_{s}$ is related to an “active variation” of the fields $\xi$ under the transformation given in Eq. (119), i.e., to a change in the fields with a fixed spacetime point. [^57] This terminology recalls, for example, the active rotation of a vector, in which the vector is changed but the coordinate system is kept fixed.
As a matter of notational convenience, we define the operator
$\delta:=s\left.\frac{\partial}{\partial s}\right|_{s=0}\!\equiv s\left(\frac{\mathrm{d}}{\mathrm{d}s}-\frac{\partial x^{\mu}}{\partial s}\frac{\partial}{\partial x^{\mu}}\right)_{s=0}\ ,$ (121)
which can be defined for any $s$, but it also yields the active variation of $\xi$ if $s$ is, in particular, very small (“infinitesimal”):
$\xi(\tau(0),\mathbf{x}(0);s)-\xi(\tau,\mathbf{x})=\delta\xi+\mathcal{O}(s^{2})\ .$ (122)
Using the Jacobian
$\det\left[\frac{\partial x^{\mu}(s)}{\partial x^{\nu}}\right]=1+s\frac{\partial}{\partial x^{\mu}}\left.\frac{\mathrm{d}}{\mathrm{d}s}\right|_{s=0}x^{\mu}(s)+\mathcal{O}(s^{2})\ ,$ (123)
and assuming that the term linear in $s$ is not identically zero, we can now compute the variation of the action to be
$S[\xi(\tau(s),\mathbf{x}(s);s)]-S[\xi(\tau,\mathbf{x})]$ (124)
$=\int\mathrm{d}^{d+1}x\,\left\{\left[\frac{\partial\mathcal{L}}{\partial\xi^{i}}-\frac{\partial}{\partial x^{\mu}}\frac{\partial\mathcal{L}}{\partial(\partial_{\mu}\xi^{i})}\right]\delta\xi^{i}\right.$
$\left.\ \ \ +\frac{\partial}{\partial x^{\mu}}\left[\frac{\partial\mathcal{L}}{\partial(\partial_{\mu}\xi^{i})}\delta\xi^{i}+s\mathcal{L}\left.\frac{\mathrm{d}}{\mathrm{d}s}\right|_{s=0}\!\!x^{\mu}(s)\right]\right\}+\mathcal{O}(s^{2})\ ,$
where $\delta\xi^{i}$ is defined according to Eq. (121). Only this “active variation” of the fields appears because the change in the fields due to the change in spacetime coordinates is canceled by the change in the integration measure [the contribution from the Jacobian factor in Eq. (123)]. The transformation is then a symmetry if it does not alter the form of the field equations, which means that the Lagrangian must change by a total time derivative:
$\int\mathrm{d}^{d}x\,\left\{\left[\frac{\partial\mathcal{L}}{\partial\xi^{i}}-\frac{\partial}{\partial x^{\mu}}\frac{\partial\mathcal{L}}{\partial(\partial_{\mu}\xi^{i})}\right]\delta\xi^{i}\right.$ (125)
$\left.\ \ \ +\frac{\partial}{\partial x^{\mu}}\left[\frac{\partial\mathcal{L}}{\partial(\partial_{\mu}\xi^{i})}\delta\xi^{i}+s\mathcal{L}\left.\frac{\mathrm{d}}{\mathrm{d}s}\right|_{s=0}\!\!x^{\mu}(s)\right]\right\}=\frac{\mathrm{d}F}{\mathrm{d}\tau}\ .$
We now assume that $\int\mathrm{d}^{d}x\,\partial_{\mathbf{x}}(\ldots)=0$ even off-shell (i.e., we restrict ourselves to field configurations that satisfy this property by means of appropriate fall-off conditions, even if they do not satisfy the field equations). Then, instead of working with $F$, we define
$G=\int\mathrm{d}^{d}x\,\left[\frac{\partial\mathcal{L}}{\partial(\partial_{\tau}\xi^{i})}\delta\xi^{i}+s\mathcal{L}\left.\frac{\mathrm{d}}{\mathrm{d}s}\right|_{s=0}\!\!\tau(s)\right]-F\ ,$ (126)
so that we can conveniently rewrite Eq. (LABEL:eq:pre-noether) as
$\int\mathrm{d}^{d}x\,\left[\frac{\partial\mathcal{L}}{\partial\xi^{i}}-\frac{\partial}{\partial x^{\mu}}\frac{\partial\mathcal{L}}{\partial(\partial_{\mu}\xi^{i})}\right]\delta\xi^{i}+\frac{\mathrm{d}G}{\mathrm{d}\tau}=0\ .$ (127)
This general result leads to both Noether theorems (see, e.g., HT:book ; Pons ; Sundermeyer ; Chataig:Thesis ; Hanson ; SalisburyRosenfeld ). Indeed, $G$ is the Noether charge, which is conserved on-shell (i.e., if the field equations are satisfied). Evidently, the precise form of $G$ depends on the Lagrangian and on the symmetry transformation considered, and it can be determined from the off-shell expression in Eq. (127) (which is valid even if the field equations are not satisfied).
As a simple example, consider the mechanical case ($d=0$) with $\xi=(q,p)$ and Lagrangian
$L=p\dot{q}-\frac{p^{2}}{2m}-V(q)\ ,$ (128)
as well as a global active time translation $\tau\mapsto\tau$, $\xi(\tau)\mapsto\xi(\tau;s)=\xi(\tau+s)$. The above results lead to $\delta\xi=s\dot{\xi}(\tau)$ which, together with Eq. (127), yields
$G=s\left[\frac{p^{2}}{2m}+V(q)\right]+\text{constant}\,.$ (129)
Since $s$ does not depend on $\tau$, the on-shell conservation of $G$ implies the on-shell conservation of the Hamiltonian (regardless of whether $s$ is very small), which is the elementary result for global time translation invariance. Evidently, we can define $G$ without any $s$ dependence by repeating the derivation of Eq. (127) using only $\partial/\partial s|_{s=0}$ instead of the $\delta$ defined in Eq. (121). This yields a charge $G$ that coincides with the Hamiltonian (up to an additive constant-in-time factor) in the case of global time translations.
For local (gauge) symmetries, the functional form of $G$ may be ascertained from the off-shell equation (127). If we adopt DeWitt’s compact notation (cf. Sec. II.1), and we take the fields to be $\xi=(z,\lambda)$ with the action given in Eq. (4) [assuming that both the symplectic potential and the Hamiltonian may depend on $z$ and its first-order spatial derivatives for consistency with Eq. (118)], then Eq. (127) becomes Eq. (13), from which we can deduce the form of $G$ and its role as the symmetry generator via the Poisson bracket [cf. Eq. (14)].